# LLM in a flash: Efficient Large Language Model Inference with Limited Memory

### v2

@@ 3.2 Improving Transfer Throughput with Increased Chunk Sizes
Figure 6: (a) For a randomly selected neuron from the 10th layer of OPT 6.7B, there exist a group of neurons which are coactivated with high probability (b) The closest friend of a neuron is defined as the most coactivated neuron in the same layer, and the closet friend of every neuron in OPT 6.7B almost always get coactivated. (c) The 3rd closest friend gets coactivatd with each neuron 86% of the time in average (d) The 7th closest friend seems to be less relevant and doesn’t coactivate with the neuron very often.
@@
We had a conjecture that neurons may be highly correlated in their activity patterns, which may enable further bundling. To verify this we calculated the activations of neurons over C4 validation dataset. For each neuron the coactivation of that neuron with other ones forms a power law distribution as depicted in Figure 5(a). Now, let’s call the neuron that coactivates with a neuron the most closest friend. Indeed, the closest friend of each neuron coactivates with it very often. As Figure 5(b) demonstrates, it is interesting to see each neuron and its closest friend coactivate with each other at least 95% of the times. The graphs for the 4th closest friend and 8th closest friend are also drawn. Based on this information we decided to put a bundle of each neuron and its closest friend in the flash memory; whenever a neuron is predicted to be active we’ll bring its closes friend too. Unfortunately, this resulted in loading highly active neurons multiple times and the bundling worked against our original intention. It means, the neurons that are very active are ‘closest friend‘ of almost everyone. We intentionally present this negative result, as we believe it may lead to interesting future research studies on how to effectively bundle the neurons and how to leverage it for efficient inference.
3.3 Optimized Data Management in DRAM
Figure 7: Memory management, first we copy last elements to deleting neurons to maintain a consecutive block of memory then the required ones are stack to the end, this prevents from copying whole data multiple times
Although data transfer within DRAM is more efficient compared to accessing flash memory, it still incurs a non-negligible cost. When introducing data for new neurons, reallocating the matrix and appending new matrices can lead to significant overhead due to the need for rewriting existing neurons data in DRAM. This is particularly costly when a substantial portion (approximately 25%) of the Feed-Forward Networks (FFNs) in DRAM needs to be rewritten. To address this issue, we adopt an alternative memory management strategy. This involves the preallocation of all necessary memory and the establishment of a corresponding data structure for efficient management. The data structure comprises elements such as pointers, matrix, bias, num_used, and last_k_active shown in Figure 7.
Each row in the matrix represents the concatenated row of the ’up project’ and the column of the ’down project’ of a neuron. The pointer vector indicates the original neuron index corresponding to each row in the matrix. The bias for the ’up project’ in the original model is represented in the corresponding bias element. The num_used parameter tracks the number of rows currently utilized in the matrix, initially set to zero. The matrix for the $i$th layer is pre-allocated with a size of $\texttt{Req}_{i}\times 2d_{\text{model}}$, where $\texttt{Req}_{i}$ denotes the maximum number of neurons required for the specified window size in a subset of C4 validation set. By allocating a sufficient amount of memory for each layer in advance, we minimize the need for frequent reallocation. Finally, the last_k_active component identifies the neurons from the original model that were most recently activated using the last $k$ tokens.
The following operations are done during inference as depicted in Figure 7.
  1. 1.
    Deleting Neurons: Neurons that are no longer required are identified efficiently in linear time, utilizing the last_k_active data and the current prediction. The matrix, pointer, and scalars of these redundant neurons are replaced with the most recent elements, and their count is subtracted from num_rows. For $O(c)$ neurons to be deleted, a memory rewrite of the order $O(c\times d_{\text{model}})$ is required.
  2. 2.
    Bringing in New Neurons: Necessary neuron data is retrieved from flash memory. The corresponding pointers and scalars are read from DRAM, and these rows are then inserted into the matrix, extending from num_row to num_row + num_new. This approach eliminates the need for reallocating memory in DRAM and copying existing data, reducing inference latency.
  3. 3.
    Inference Process: For the inference operation, the first half of the matrix[:num_rows,:d_model] is used as the ’up project’, and the transposed second half, matrix[:num_rows,d_model:].transpose(), serves as the ’down project’. This configuration is possible because the order of neurons in the intermediate output of the feed-forward layer does not alter the final output, allowing for a streamlined inference process.
4 Results
Experimental Setup: Our experiment is designed to optimize inference efficiency on personal devices. To this end, we process sequences individually, running only one sequence at a time. This approach allows us to allocate a specific portion of DRAM for the Key-Value (KV) cache while primarily focusing on the model size. This strategy is particularly effective when dealing with only one sequence/query at a time.222 For OPT 6.7 B model with context length 2048 KV-cache requires $2048\times 2d_{\text{model}}$ elements which is only $8\%$ of model size. Also the KV-cache itself can be held in flash memory.
For the implementation of our inference process, we utilize the HuggingFace’s transformers and KV caching. This setup is tested under the condition where approximately half of the model size is available in DRAM. We select this amount as a showcase of the idea of hosting the LLM in flash. With a different level of sparsity or employing quantization, one can work with smaller available DRAM capacity as well. Such a configuration demonstrates the practicality of executing inference with lower memory footprints.
Hardware Configuration. Our models are evaluated using two distinct hardware setups. The first setup includes an Apple M1 Max with a 1TB solid-state drive (SSD) for flash memory. In this configuration, computations are performed on the CPU, and the models are maintained in a 32-bit format. The second setup involves a Linux machine equipped with a 24 GB NVIDIA GeForce RTX 4090 graphics card. For this machine, computations are GPU-based, and models are run in the bfloat16 format. For both setups, we operate under the assumption that almost half of the total available memory (DRAM plus GPU memory) is allocated for model computations.
Models. We use OPT 6.7B (Zhang et al., 2022b) and a sparsified Falcon 7B (Mirzadeh et al., 2023) model for our evaluations.
Baselines. For methods not employing sparsity or weight sharing, at least half of the model must be transferred from flash memory during the forward pass. This necessity arises because, initially, only half of the model is available in DRAM, but as the forward pass progresses, the entire model capacity is utilized. Consequently, any data not present at the start must be transferred at least once. Thus, the most efficient theoretical baseline involves loading half of the model size from the flash memory into DRAM. This optimal I/O scenario serves as our primary baseline. Comparative methods, such as FlexGen (Sheng et al., 2023) and Petals (Borzunov et al., 2023), are also constrained by the limited available DRAM or GPU memory, and therefore cannot surpass this theoretical I/O efficiency.
Flash memory Data Loading Implementation. To optimize data loading from flash memory, our system employs reads parallelized over 32 threads. This multithreaded approach is intended to both better amortize latency to first byte by not waiting for each read sequentially, and maximize read throughput by reading multiple streams at once (Figure 1(b)).
Caching Considerations for Data Loading from Flash Memory. When data is read from flash memory, the operating system typically caches these pages, anticipating future reuse. However, this caching mechanism consumes additional memory in DRAM beyond what is allocated for the model. To accurately assess the real throughput of flash memory under limited DRAM conditions, benchmarks should be conducted without relying on caching. Practical systems may or may not rely on filesystem cache, depending on requirements.
For the purpose of our hardware benchmarking in this study, we deliberately and significantly pessimize our NVMe throughput measurements. On macOS and iOS, we employ the F_NOCACHE flag with the fcntl() function, while on Linux, we use DirectIO. Additionally, on macOS, we clear any resident buffers before initiating the benchmark using the purge command. This approach provides a conservative lower bound of throughput in scenarios where no caching is permitted, and makes the benchmarks repeatable. It’s worth noting that these figures can improve if either the inference code or the operating system is allowed to cache some part of the weights.
While OS-level buffer caching is advantageous for general purpose applications with high cache hit rates, it lacks fine-grained control over cache usage per process or buffer eviction at the application level. In the context of on-device memory constraints and large model sizes, this could lead to a situation where filesystem level does not help, because in order to evaluate later layers earlier layers must be evicted in a rolling pattern, so the effective cache hit rate is close to zero. Aside from being inefficient, this can cause coexistence issues with other processes due to memory allocation pressure and Translation Lookaside Buffer (TLB) churn.
4.1 Results for OPT 6.7B Model
Predictors. For the initial 28 layers of the OPT 6.7B model, we train predictors with a rank of $r=128$. To reduce the occurrence of false negatives, the final four layers employ predictors with a higher rank of $r=1024$. These predictors achieve an average of 5% false negatives and 7% false positives in the OPT 6.7B model. As depicted in Figure 2(a), our predictor accurately identifies most activated neurons, while occasionally misidentifying inactive ones with values near zero. Notably, these false negatives, being close to zero, do not significantly alter the final output when they are excluded. Furthermore, as demonstrated in Table 1, this level of prediction accuracy does not adversely affect the model’s performance in 0-shot tasks.
Windowing in the OPT 6.7B Model. Utilizing a windowing method with $k=5$ in the OPT 6.7B model significantly reduces the necessity for fresh data loading. Using active neurons of predictor would require about 10% of the DRAM memory capacity in average; however, with our method, it drops to 2.4%. This process involves reserving DRAM memory for a window of the past 5 tokens, which, in turn, increases the DRAM requirement for the Feed Forward Network (FFN) to 24%.
Table 2: The I/O latency of OPT 6.7B 16 bit on M1 max for different techniques when half the memory is available
Configuration Performance Metrics
Hybrid Predictor Windowing Bundling DRAM (GB) Flashnormal-$\rightarrow$ DRAM(GB) Throughput (GB/s) I/O Latency (ms)
0 13.4 GB 6.10 GB/s 2130 ms
6.7 6.7 GB 6.10 GB/s 1090 ms
4.8 0.9 GB 1.25 GB/s 738 ms
6.5 0.2 GB 1.25 GB/s 164 ms
6.5 0.2 GB 2.25 GB/s 87 ms
Latency Analysis: Using a window size of 5, each token requires access to 2.4% of the Feed Forward Network (FFN) neurons. For a 32-bit model, the data chunk size per read is $2d_{\text{model}}\times 4\text{ bytes}=32\text{ KiB}$, as it involves concatenated rows and columns. On an M1 Max, this results in the average latency of $125ms$ per token for loading from flash and $65ms$ for memory management (involving neuron deletion and addition). Thus, the total memory-related latency is less than $190ms$ per token (refer to Figure 1). In contrast, the baseline approach, which requires loading 13.4GB of data at a speed of 6.1GB/s, leads to a latency of approximately $2330ms$ per token. Therefore, our method represents a substantial improvement over the baseline.
For a 16-bit model on a GPU machine, the flash load time is reduced to $40.5ms$, and memory management takes $40ms$, slightly higher due to the additional overhead of transferring data from CPU to GPU. Nevertheless, the baseline method’s I/O time remains above 2000 milliseconds.
4.2 Results for Falcon 7B Model
To verify that our findings generalize beyond OPT models we also apply the idea of LLM in flash to Falcon model. Since, the base line Falcon model is not sparse, we used a sparsified (relufied) version with almost the same performance as that of the base version (Mirzadeh et al., 2023). Similar to previous section, we present the results obtained under the condition that approximately half of the model size is available for use in DRAM.
Latency Analysis. Using a window size of 4 in our model requires accessing 3.1% of the Feed Forward Network (FFN) neurons for each token. In a 32-bit model, this equates to a data chunk size of 35.5 KiB per read (calculated as $2d_{\text{model}}\times 4\text{ bytes}$). On an M1 Max device, the time taken to load this data from flash memory is approximately 161ms, and the memory management process adds another 90ms, leading to a total latency of 250ms per token. In comparison, the baseline latency is around 2330 milliseconds, making our method approximately 9 to 10 times faster.
5 Related Works
@@
Efficient Inference for Large Language Models. As LLMs grow in size, reducing their computational and memory requirements for inference has become an active area of research. Approaches broadly fall into two categories: model compression techniques like pruning and quantization (Han et al., 2016b; Sun et al., 2023; Jaiswal et al., 2023; Xia et al., 2023), (Zhang et al., 2022a; Xu et al., 2023; Shao et al., 2023; Lin et al., 2023; Hoang et al., 2023; Zhao et al., 2023; Ahmadian et al., 2023; Liu et al., 2023a; Li et al., 2023), and selective execution like sparse activations (Liu et al., 2023b), (Mirzadeh et al., 2023) or conditional computation (Graves, 2016; Baykal et al., 2023). Our work is complementary, focusing on minimizing data transfer from flash memory during inference.
@@
Selective Weight Loading. Most related to our approach is prior work on selective weight loading. SparseGPU (Narang et al., 2021) exploits activation sparsity to load a subset of weights for each layer. However, it still requires loading from RAM. Flexgen (Sheng et al., 2023) offloads the weights and kv-cache from GPU memory to DRAM and DRAM to flash memory, in contrast we consider only the cases the full model can’t reside in the whole DRAM and GPU memory on the edge devices. Flexgen is theoretically bound by the slow throughput of flash to DRAM in such scenarios. Firefly (Narang et al., 2022) shares our goal of direct flash access but relies on a hand-designed schedule for loading. In contrast, we propose a cost model to optimize weight loading. Similar techniques have been explored for CNNs (Parashar et al., 2017), (Rhu et al., 2013). Concurrently, Adapt (Subramani et al., 2022) has proposed adaptive weight loading for vision transformers. We focus on transformer-based LLMs and introduce techniques like neuron bundling tailored to LLMs.
@@
Hardware Optimizations. There is a rich body of work on hardware optimizations for efficient LLM inference, including efficient memory architectures (Agrawal et al., 2022), (Gao et al., 2022), dataflow optimizations (Han et al., 2016a), (Shao et al., 2022), hardware evaluation frameworks Zhang2023AHE, and flash optimizations (Ham et al., 2016), (Meswani et al., 2015). We focus on algorithmic improvements, but these could provide additional speedups.
Speculative Execution. Speculative decoding (Leviathan et al., 2022; Zhang et al., 2023; He et al., 2023) is a technique that uses a draft model for generation and uses the larger model to verify those tokens. This technique is orthogonal to us and can be used for further improvement. In case of speculative decoding, the window in our method should be updated with multiple tokens rather one.
Mixture of Experts. Mixture of Experts (Yi et al., 2023) have a sparse structure in their feed forward layer and can leverage our method for enabling larger models on device.
@@
6 Conclusion and Discussion
In this study, we have tackled the significant challenge of running large language models (LLMs) on devices with constrained memory capacities. Our approach, deeply rooted in the understanding of flash memory and DRAM characteristics, represents a novel convergence of hardware-aware strategies and machine learning. By developing an inference cost model that aligns with these hardware constraints, we have introduced two innovative techniques: ’windowing’ and ’row-column bundling.’ These methods collectively contribute to a significant reduction in the data load and an increase in the efficiency of memory usage. Weight bundling and windowing are two very basic techniques aimed at showcasing the potentials to increase chunk size and read sequentiality while reducing data transfer through sparsity. Numerous opportunities exist for developing smarter and more efficient methods to achieve these objectives.
The practical outcomes of our research are noteworthy. We have demonstrated the ability to run LLMs up to twice the size of available DRAM, achieving an acceleration in inference speed by 4-5x compared to traditional loading methods in CPU, and 20-25x in GPU. This innovation is particularly crucial for deploying advanced LLMs in resource-limited environments, thereby expanding their applicability and accessibility. The PyTorch based implementation for forward pass have only undergone algorithmic (as opposed to systems) optimization. Significant additional gains are expected from a custom lower level implementation.
Our work not only provides a solution to a current computational bottleneck but also sets a precedent for future research. It underscores the importance of considering hardware characteristics in the development of inference-optimized algorithms, suggesting a promising direction for further explorations in this domain. We believe as LLMs continue to grow in size and complexity, approaches like this work will be essential for harnessing their full potential in a wide range of devices and applications.
Our study represents an initial endeavor in the pursuit of democratizing Large Language Model (LLM) inference, making it accessible to a wider array of individuals and devices. We recognize that this early effort has its limitations, which, in turn, open up compelling avenues for future research. A critical aspect for future exploration is the analysis of power consumption and thermal limitations inherent in the methods we propose, particularly for on-device deployment. Currently, our focus is on single-batch inference. However, expanding this to include scenarios like prompt processing, multi-batch inference, and speculative decoding presents itself as a valuable area for further investigation. In our initial proof of concept, we operated under the assumption of memory availability being half the size of the model. Exploring the dynamics of working with varying memory sizes—both larger and smaller—introduces a fascinating balance between latency and accuracy, and is a compelling area for future exploration. In conclusion, our methodology is constructed on the foundation of sparsified networks. Nonetheless, the underlying concept holds potential for broader applications. It can be adapted to selectively load weights in non-sparse networks or to dynamically retrieve model weights from flash storage. This adaptation would be contingent on the specific requirements of the input prompt or the contextual parameters provided. Such an approach suggests a versatile strategy for managing model weights, optimizing performance based on the nature of the input, thereby enhancing the efficiency, usefulness, and applicability of the proposed scheme in various scenarios dealing with Large Language Models (LLMs).
Acknowledgements
We would like to thank Itay Sagron, Lailin Chen, Mahyar Najibi, Qichen Fu, Moin Nabi, Peter Zatloukal, Arsalan Farooq, Sachin Mehta, Mohammad Samragh, Matt Johnson, Etai Zaltsman, Lin Chang, Dominic Giampaolo, Taal Uliel, Hadi Pouransari, Fartash Faghri, Oncel Tuzel, Samy Bengio, Ruoming Pang, Chong Wang, Ronan Collobert, David Grangier, and Aftab Munshi for the valuable feedback and discussions.

### v3

@@ Appendix D Bundling Based on Co-activation
Figure 12: (a) For a randomly selected neuron from the 10th layer of OPT 6.7B, there exists a group of neurons that are coactivated with high probability (b) The closest friend of a neuron is defined as the most coactivated neuron in the same layer, and the closet friend of every neuron in OPT 6.7B almost always get coactivated. (c) The 3rd closest friend gets co-activated with each neuron 86% of the time on average (d) The 7th closest friend seems to be less relevant and doesn’t coactivate with the neuron very often.
@@
Given the high reuse of data in sparse models, we hypothesize that neurons may be highly correlated in their activity patterns, which may enable further bundling. To verify this we calculated the activations of neurons over the C4 validation dataset. For each neuron, the coactivation of that neuron with other ones forms a power law distribution as depicted in Figure  11(a). Now, let’s call the neuron that coactivates with a neuron the most closest friend. Indeed, the closest friend of each neuron coactivates with it very often. As Figure  11(b) demonstrates, it is interesting to see each neuron and its closest friend coactivate with each other at least 95% of the time. The graphs for the 4th closest friend and 8th closest friend are also drawn. Based on this information we decided to put a bundle of each neuron and its closest friend in the flash memory; whenever a neuron is predicted to be active we’ll bring its closest friend too. Unfortunately, this resulted in loading highly active neurons multiple times and the bundling worked against our original intention. It means the neurons that are very active are the ‘closest friends’ of almost everyone.
Appendix E Extended Related Works
@@
Efficient Inference for Large Language Models. As LLMs grow in size, reducing their computational and memory requirements for inference has become an active area of research. Approaches broadly fall into two categories: model compression techniques like pruning and quantization (Han et al., 2016b; Sun et al., 2023; Jaiswal et al., 2023; Xia et al., 2023), (Zhang et al., 2022a; Xu et al., 2023; Shao et al., 2023; Lin et al., 2023; Hoang et al., 2023; Zhao et al., 2023; Ahmadian et al., 2023; Liu et al., 2023a; Li et al., 2023), and selective execution like sparse activations (Liu et al., 2023b; Mirzadeh et al., 2023; Zhang et al., 2024) or conditional computation (Graves, 2016; Baykal et al., 2023). Our work is complementary, focusing on minimizing data transfer from flash memory during inference.
@@
Selective Weight Loading. Most related to our approach is prior work on selective weight loading. Dejavu (Liu et al., 2023b) exploits activation sparsity to load a subset of weights for each layer. However, it still requires loading from GPU memory. Flexgen (Sheng et al., 2023) offloads the weights and KV-cache from GPU memory to DRAM and DRAM to flash memory, in contrast, we consider only the cases where the full model can’t reside in the whole DRAM and GPU memory on the edge devices. Flexgen is theoretically bound by the slow throughput of flash to DRAM in such scenarios. Similar techniques have been explored for CNNs (Parashar et al., 2017), (Rhu et al., 2013). Concurrently, Adapt (Subramani et al., 2022) has proposed adaptive weight loading for vision transformers. We focus on transformer-based LLMs and introduce techniques like neuron bundling tailored to LLMs.
@@
Hardware Optimizations. There is a rich body of work on hardware optimizations for efficient LLM inference, including efficient memory architectures (Gao et al., 2022), dataflow optimizations (Han et al., 2016a; Shao et al., 2022), hardware evaluation frameworks (Zhang et al., 2023a), faster sparse kernels (Gale et al., 2020) and flash optimizations (Ham et al., 2016), (Meswani et al., 2015). We focus on algorithmic improvements, but these could provide additional speedups.
Speculative Execution. Speculative decoding (Leviathan et al., 2022; Zhang et al., 2023b; He et al., 2023) is a technique that uses a draft model for generation and uses the larger model to verify those tokens. This technique is orthogonal to us and can be used for further improvement. In the case of speculative decoding, the window in our method is updated with multiple tokens rather than one.
Mixture of Experts. Mixture of Experts (Yi et al., 2023) have a sparse structure in their feed-forward layer and can leverage our method for enabling larger models on the device.
@@
Appendix F Small Device Implications
Table 6: Active neuron percentage in different layers of OPT 6.7B vs Quantized model over 100 sequences.
Layer OPT 6.7B Quantized
1 1.56% 1.42%
16 2.66% 2.44%
32 5.36% 5.45%
[…]
We note that many of the hardware assumptions (e.g., limited DRAM capacity, characteristics of Flash such as bandwidth limitations and increased throughput with larger chunks) are also applicable to smaller devices like smartphones. For example, when running a 7B model on a smartphone, the same technique can be employed; instead of loading 16-bit values, 4-bit values would be loaded. Consequently, rather than the baseline’s 3.5GB of required DRAM, our model would necessitate less than 2GB of memory for operation.x We believe that since quantization does not alter activation sparsity patterns, the same technique can be utilized here by loading 4-bit instead of 16-bit values (Table 6). However, actual implementation would require special 4-bit compute kernels on device, which falls outside the scope of this paper.
Appendix G Qualitative Evaluations
Here we present some of the examples generated from the model with a predictor. The purpose is to qualitatively demonstrate that the models’ output is comparatively reasonable when the predictors are applied. Table 7 and 8 show the results where the prompt was "once upon a time there was a" and the models are asked to complete it. Table 9 shows the results after prompting the model to complete a random sample taken from C4 dataset.
Table 7: Comparison of the output of original model and model with predictors.
Prompt: Once upon a time there was a
OPT 6.7B OPT 6.7B with predictors
little girl who was very sad. She was sad because her daddy was gone. He was in the army and he was in Iraq. She was sad because she missed him so much. She was sad because she was so lonely. She was sad because she was so scared. She was sad because she was so sad. She was sad because she was sad. She was sad because she was sad. She was sad because she was sad. man who was very rich. He had a beautiful wife and two beautiful children. He was a very good man and he loved his family very much. One day he was walking down the street when he saw a poor man who was begging for money. The man was very poor and he had no money. The man asked the man for some money. The man gave the man some money. The man was very happy and he thanked the man.
Persimmon 8B Persimmon 8B with predictors
man named Bilbo Baggins. Bilbo Baggins was a small hobbit, and he lived in a hole in the ground. Bilbo Baggins had many adventures, and he met many friends. Bilbo Baggins was brave, and he was not afraid of anything. Bilbo Baggins was happy, and he always smiled. Bilbo Baggins was a kind person, and he always helped others. man named Bilbo Baggins. He lived in a hole in the ground, with many other hobbits. Bilbo Baggins was a very brave hobbit. He liked to eat LOTS of cheese. One day, Bilbo Baggins decided to leave his hobbit hole. He decided to travel the world. Bilbo Baggins was very hungry. He ate many cheeses. Bilbo Baggins was very tired. He slept in a cave. Bilbo Baggins was very cold. He
Table 8: Comparison of the output of original model and model with predictors.
Prompt: Once upon a time there was a
Phi-2 Relufied Phi-2 with predictors
young girl named Lily who loved to read. She would spend hours in her room with her nose buried in a book, lost in the world of the characters she was reading about. Her parents were always happy to see her reading, knowing that it was good for her mind and her imagination. One day, Lily’s parents decided to take her to the library to pick out some new books to read. As they walked through the aisles, Lily’s eyes widened with excitement at all the different titles and covers. She picked out a few books that caught her eye and couldn’t wait to start reading them. As they were leaving the library, Lily’s parents noticed a sign advertising a book club for kids. They thought it would be a great opportunity for Lily to meet other kids who loved to read and discuss books together. Lily was hesitant at first, but her parents convinced her to give it a try. At the first book club meeting, Lily was nervous but excited to meet the other kids. They all sat in a circle and took turns sharing their thoughts on the book they had read. Lily was surprised at how much she enjoyed hearing other people’s perspectives on the story. She realized that reading wasn’t just about escaping into a world of her young girl named Lily who loved to read. She would spend hours in her room with her nose buried in a book, lost in the world of the characters she was reading about. Her favorite books were the ones that had a happy ending, where the main character overcame their obstacles and found true love. One day, Lily’s parents surprised her with a trip to the library. She was so excited to see all the new books she could check out. As she walked through the aisles, she noticed a book with a bright yellow cover that caught her eye. It was called "The Adventures of Lily and the Magic Book." Lily eagerly checked out the book and took it home to read. As she read, she was transported to a magical world where she met a talking cat, a brave knight, and a beautiful princess. The book was filled with adventure, magic, and romance. As Lily read on, she noticed that the book was getting darker and darker. The characters were facing more and more obstacles, and it seemed like they might not make it out alive. Lily began to worry that the book was going to end badly.
Sparse Llama 2 Sparse Llama 2 with predictors
little girl who loved to play with her dolls. She loved to dress them up and make them do all sorts of things. One day, she decided to make a dollhouse for her dolls. Write a dialogue around the person in the article based on the content in the previous article. Melanie: Hey, Mom! I was thinking about making a dollhouse for my dolls. Mother: That sounds like a great idea, Melanie! What little girl named Lily. She loved to play with her dolls and stuffed animals. One day, she decided to make a new doll for herself. Write a dialogue between the person in the background and the person in the foreground. Person in the background: Lily, what are you doing? Person in the foreground: I’m making a doll for myself. Person in the background: Why do you need a doll? Person in the foreground
Table 9: Comparison of the output of original model and model with predictors.
Prompt: There is a lot of discussion right now about stress reduction. Today we see an increased work load of mental health professionals, increased occurrences of abuse or violence faced by social workers and other therapists, and the terrible struggles facing our returning veterans. All of these things have increased our awareness of the stressful elements of our society and the impact those elements have on both our clients and the professionals whose job it is to help. As a social worker I am
Persimmon 8B (nucleus sampling) Persimmon 8B with predictors (nucleus sampling)
constantly reminded of the importance of stress reduction. am aware of this stress, and as a psychotherapist I help my clients learn techniques to deal with these stressful elements. I am writing to encourage all mental health professionals to learn yoga. I strongly believe that yoga should be part of your clinical routine, and that yoga classes should be taken by clinicians for credit, just like psychotherapy, case management and substance use classes constantly reflecting on my own stress levels and how I can remain more fit to handle the pressures of the profession. In social work a social worker calls themselves "a problem solver." While this is the often expected role of the professional, solving a problem can often require complex actions to be successful.
Phi-2 Relufied Phi-2 with predictors
constantly looking for ways to reduce stress in my life and in the lives of my clients. One of the most effective ways to reduce stress is to take time to relax. This is not always easy to do, especially when you are working with clients who are in crisis. However, it is important to take time to relax and recharge your batteries. One way to do this is to take a break from work and go for a walk. Walking is a great way to clear your mind and get some often asked how I deal with stress. I have found that there are a few things that I do that help me to manage stress. 1. I make sure that I get enough sleep. I try to get at least 8 hours of sleep a night. I also try to get up early and get out of bed by 7:00 a.m. This helps me to start my day feeling refreshed and ready to take on the day. 2. I make sure that I eat a healthy diet